Could someone please help me, Thank You so much.
Is x = 6 a solution to -2x – 4 = -8?

Answer by nerdybill(7384)

Is x = 6 a solution to -2x – 4 = -8?
To check replace the 'x' with a 6 in:
-2x – 4 = -8
-2(6) – 4 = -8
-12 – 4 = -8
-16 = -8 NOT true
.
Since it's not true, x=6 is NOT a solution